OpenStudy (anonymous):

Antiderivative help f ''(t) = 9/t^(1/2), f(4) = 32, f '(4) = 9 find f(t)

OpenStudy (anonymous):

\[f''(t)=\frac{ 9 }{ \sqrt{t} } \therefore f''(t)=9t ^{-1/2}\]

OpenStudy (shamil98):

f''(t) = 9 / t^(1/2) = 9 * t^(-1/2) can you integrate it from here?

OpenStudy (anonymous):

the problem im having is with the constants. i am getting \[12t ^{2/3}+t/4-65 \] after integrating twice. if it were not for the -65 i would be correct

OpenStudy (anonymous):

what did you get before that step?

OpenStudy (anonymous):

\[18\sqrt{t}+t/4\]

OpenStudy (anonymous):

okay thats good, factor out a t^1/2

OpenStudy (shamil98):

omg, i had the latex written and accidentally deleted

OpenStudy (shamil98):

:( , hold on i'll rewrite it >.>

OpenStudy (anonymous):

Thats the worst

OpenStudy (anonymous):

how do you factor t^1/2 out of t/4

OpenStudy (anonymous):

also it should be \[18\sqrt{t}+t/4+c\]

OpenStudy (anonymous):

since\[f'(x)=18\sqrt{t}+t/4+c\] and f'(4)=9 you should get c= 1/4

OpenStudy (anonymous):

alright, so i got a answer : \[12t^\frac{ 3 }{ 2 }-27t+92...\]

OpenStudy (anonymous):

i believe

OpenStudy (anonymous):

t^1/2 [ 18+( (t^1/2)/4) ] would be how you factor it

OpenStudy (anonymous):

alanMW web assign says your wrong too. :( the amount of times ive worked this out i swear i should have gotten it right

OpenStudy (anonymous):

ive given up. i ran out of guesses. :( thanks anyways though everyone! if you find an answer let me know!

OpenStudy (anonymous):

actually i got \[12t^\frac{ 3 }{ 2 }-27t+44\]

OpenStudy (anonymous):

i am about to close this and ask another question about antiderivatives. @AlanMW i cant check it or i would tell you if you are right.

OpenStudy (anonymous):

im thinking thats it

OpenStudy (anonymous):

@alanmw if you take the derivative of it twice it comes out right, but my answer did as well. ill give you a medal though

OpenStudy (anonymous):

It can be factored like this (sorry i forgot to push send) \[18\sqrt{t} + \frac{ t }{ 4 } ------------>>>>>>> 2 t ^{\frac{ 1 }{ 2 }} [ 9 + \frac{ ^{t ^{\frac{ 1 }{ 2 }}} }{ 8 }]\]